Sam can paint the fence in 9 hours and Bob can paint it in 6 hours. How long will it take if Sam and Bob paint the fence together?
I tried 1/9 + 1/6 = 1/15 so I got the answer of "3 hours" but it was marked wrong.

Answer by rwm(914)

that's because it is wrong.
1/9+1/6= 6/54+9/54=15/54=5/18 not 1/15
how on earth did you get an answer of three from
1/9+1/6=1/15 there is no variable?
You just through a bunch of numbers on your paper.

Together they do one -1- job
x is the number of hours they work
x/9+x/6=1
multiply by 18
2x+3x=18
5x=18
x=18/5 =3 3/5 hrs